At a baseball game, the batter hits the baseball straight up at an initial velocity of +42m/s.
A) at what time will the ball reach its highest point
b) what will be the maximum height the ball reaches
BTW the answer is about 4.3 secs for part a and 90 m for part b.. i dont know how to solve it --- would like to see your steps. THANKS!!

Answer by Alan3354(69443)

At a baseball game, the batter hits the baseball straight up at an initial velocity of +42m/s.
--------------------
h(t) = gt^2/2
Using 10 m/sec/sec for gravity:
h(t) = -5t^2 + 42t
That's a parabola.
The max height is the vertex, at t = -b/2a
t = -42/-10
Using 9.8 m/sec/sec gives ~ 4.28 seconds
============================
Max height is at t = 4.28
h(2.14) = -4.9*4.28^2 + 42*4.28
= ~90 meters

A) at what time will the ball reach its highest point
b) what will be the maximum height the ball reaches